NASA ISS On-Orbit Status 18 December 2013


NASA ISS On-Orbit Status 18 December 2013

Today: External Thermal Control System (ETCS) Loop A Status: Following yesterday's S1-1 MDM software patch, Flight Controllers continued to perform characterization tests on ETCS Loop A to regain fine temperature control of ETCS loop A.

The Node 2 Moderate Temperature Loop (MTL) IFHX is still isolated from ETCS Loop A, resulting in half of the loads in the Node 2, JEM and Columbus modules remaining powered off.

Extravehicular Activity (EVA) Mobility Unit (EMU) 3011 Preparation: This morning, Flight Engineer (FE)-3 Hopkins and FE-5 Mastracchio removed the medium sized HUT on EMU 3011 and replaced it with a large sized HUT. In addition, they swapped out the Secondary Oxygen Packs (SOP) on the EMU. Later in their day, Mastracchio verified that the reassembled EMU was fully functional. USOS EVA's are scheduled to occur on December 21st, 23rd, and 25th (if required) in order to repair the ETCS Loop A Pump Module.

Ocular Health (OH) Optical Coherence Tomography (OCT) Test and Fundoscope Examination Deferred: Today's OH and OCT sessions were deferred in order to allow enough time to support EVA related activities.

Today's Planned Tasks

Hard Upper Torso (HUT) changeout on EMU 3011 [Complete]
Microbial Removal Filter (MRF) - Filter Purge [Complete]
Activation and Inspection of Orlan Suits No.4, No. 6 and Checkout of Spacesuit Interface Unit [Complete]
SM Ventilation System Preventive Maintenance. Group В1 [Complete]
CMS Battery Replacement [Complete]
Orlan Suit No. 4, No. 6, and БСС Cooling Loop Separation in DC1 [Complete]
Profile of Mood States (POMS) Questionnaire [Complete]
Hydrogen (H2) Sensor ORU R&R, and Air Avionics Assembly (AAA) Cleaning Part 2 of 2 [Complete]
Orlan Replaceable Elements Ops, Orlan Inspection [Complete]
Mating connectors of RACU 5, 6. [Complete]
Microbial Removal Filter (MRF) - Filter Purge [Complete]
Spacesuit and БСС (Orlan Interface Unit) Leak Check and Orlan Valve Functional Tests [Complete]
Installation of EMU 3011 upside down in EDDA in preparation for SOP R&R. [Complete]
Journals Nominal Entry [Complete]
Secondary Oxygen Packs (SOP) between two EMUs and perform checkout [Complete]
Terminate 825М3 Battery Pack 1 Charging and Start Charging 825М3 Battery Pack 2 [Complete]
ACS High Pressure Oxygen Bypass Configuration [Complete]
EMU 3011 Resize [Complete]
Terminate transfer of water to the water storage tank via the MRF. [Complete]
CB/Crew Conference [Complete]
EMU Checkout [Complete]
Optical Coherence Tomography (OCT) Test Setup [Complete]
Equipment Setup [Complete]
Optical Coherence Tomography (OCT) Test [Complete]
Filling EDV-SV with water from ТКГ 421 (AO) Rodnik Tank 1 [Complete]
БСС Water Loop Separation in ПхО Using Degassing Pump Unit (БОС) [Complete]
Air Heater Fan (БВН) Screen Cleaning in ТК 710 [Complete]
СОЖ Maintenance [Complete]
Optical Coherence Tomography (OCT) Test Stow [Complete]
EMU Swap [Complete]
WRM Condensate Swap [Complete]
Fundoscope Setup [Complete]
Fundoscope Prep [Complete]
Collect SM and FGB Air Samples Using АК-1М Sampler [Complete]
SM Air Sampling for CO using ИПД Samplers [Complete]
IMS Update [Complete]
Fundoscope Exam [Complete]
Fundoscope Stow [Complete]
Task List

ARED Cylinder Evacuation
ТКГ 421 Transfers and IMS Updates /Transfer Ops
Replace KTO when full.
CWC-I (Iodinated Contingency Water Container) Degas.
Preparation of reports for Roskosmos site
URAGAN. Observations and Photography
ECON-M. Observations and Photography
Prints new cuff check list pages for EVAs [In Work]
Procedure Review
National Geographic B-Role
OGS Jumper Routing
P/TV EVA camera Config
PAO Event
IFM stowage Locate
Kubik 6 Belt Search
National Engr Week Message
P/TV Trash
PPS Wipe
Pre-gather COL A/L
Pre-gather Empty CTBs
SM PCS HD Swap Imaging

Nominal Commanding [Complete]
PMCU MDM Swap [Complete]
Three-Day Look Ahead:

Thursday, 12/19: PHS, EVA decontamination training, EVA Procedure Review, Vision Test
Friday, 12/20: EVA Tool Config, EVA Camera Config, Equipment Lock Prep, EVA Procedure Review
Saturday, 12/21: USOS EVA 24
QUICK ISS Status - Environmental Control Group:

Component - Status
Elektron - On
Vozdukh - Manual
[СКВ] 1 - SM Air Conditioner System ("SKV1") - On
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Off
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) - Standby
Urine Processing Assembly (UPA) - Standby
Trace Contaminant Control System (TCCS) Lab - Off
Trace Contaminant Control System (TCCS) Node 3 - Full up

Please follow SpaceRef on Twitter and Like us on Facebook.